OpenStudy (anonymous):

Can someone help me figure this out?? Factor the expression -4x^2+14x-6

OpenStudy (tkhunny):

First, \(-2(2x^{2} - 7x + 3)\) There are very few possibilities from here. List them!

OpenStudy (anonymous):

i was told i need to take 3*2 and get 6 then i would get 2x^2 -1x -6x +3

OpenStudy (anonymous):

when i go an more i get lost

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

\[-2(2x^2-7x+3)=-2(2x-a)(x-b)\] and \(ab=3\) you have only two choices

OpenStudy (anonymous):

\[-2(2x-3)(x-1)\] or \[-2(2x-1)(x-3)\] pick the one that works

OpenStudy (anonymous):

not sure im lost, the way i was told to do it work for the last few but not this one something called the xfactor where you take the 3*2 to get 6 then take the -7 find some thing that you can multiply to get 6 but add to get -7

OpenStudy (anonymous):

ok thank you this one got me

OpenStudy (anonymous):

I figured it out the other way now that i had the answer. -4x^2+14x-6 -2(2x^2-7x+3) 3*2=6 -1*-6=6 and -1+-6=-7 sub -7x with -1x and -6x (2x^2-x)+(-6x+3) x(2x-1)-3(2x-1) so the answer would be -2(2x-1)(x-3)